Find the equation of the tangent line to the curve y=2sec(x)−4cos(x) at the point (π/3,2). Write your answer in the form y=mx+b where m is the slope and b is the y-intercept.
1
Expert's answer
2020-03-20T18:35:59-0400

Slope = m = dy/dx , at point (π /3 ,2)

m = 2 secx tanx + 4 sinx ,at point(π /3 ,2)

m = 6✓3

Y = mX + b

Now we will find value of m at point (π/3 , 2)

b = Y - mX = 2 - 6√3 * π/3 = -8.88

Y = 10.39X - 8.88